Find f ’(x) if
f(x) = 8ln(x) - 7cos(x) + 1
and evaluate f ’(1).(Keep 2 decimal places)

Answer by fractalier(6550)

For
f(x) = 8ln(x) - 7cos(x) + 1
f'(x) = 8/x + 7sin(x)
f'(1) = 8 + 5.89 = 13.89